# MASA: Rethinking the Representational Bottleneck in LoRA with Multi-A Shared Adaptation

### v1

@@ Appendix C Appendix B Rationale for Summation-Based Aggregation
Token-level routers, introduced by the Sparsely-Gated layer and subsequently adopted in GShard and Switch-Transformer, enable conditional computation (top-$k$ expert selection) and explicit expert specialization. These benefits explain their popularity in large-scale language and vision models.
@@
  • SmoothMoE replaces top-$k$ gating with a soft mask. With identical expert backends, the dense variant stays within $0.1$ BLEU of the routed model on WMT En-De while reducing gradient variance and eliminating auxiliary losses.
  • On ImageNet, V-MoE, reports that a dense (all-expert) mixture narrows the accuracy gap with routed ViT-MoE to less than $0.3$ pp, while avoiding load-balancing issues.
  • UnifiedMoE shows that additive expert fusion is at most $0.2$ bits / char behind Switch-Transformer on C4, yet is considerably easier to train.
@@
The proposed architecture prioritizes feature-side parameter enrichment through multi-$A$ experts while implementing asymmetric cross-layer sharing (ASC) to achieve parameter efficiency. The uniform summation mechanism, expressed as $\Delta W_{\text{MAE}}=B\left(\sum_{i=1}^{N}A_{i}\right)$, provides three key advantages:
@@
    Empirical sufficiency: While sparse routing strategies combined with effective load balancing have been proven to facilitate expert differentiation, our visualization results reveal that even a simple summation strategy can achieve effective expert specialization. The experimental results presented in Figure2 (Section3) provide clear evidence of task differentiation among our Expert A matrices, demonstrating that summation mechanism is sufficient for achieving functional specialization.
@@
Appendix D Appendix C Experimental Setting
Datasets and Benchmarks
@@ Appendix D Appendix C Experimental Setting
Model MMLU Params. (%) AA $B$
Hum. STEM Social Other Avg.
LLaMA3.2-3B 24.36 21.15 21.90 23.98 23.02
Prompt Tuning 25.36 24.52 24.50 23.95 25.36 0.0008
P-Tuning 25.08 25.06 21.81 24.46 24.44 0.0530
LoRAr=8 26.14 24.26 25.58 28.81 26.19 0.3770 1 1
LoRAr=16 29.73 28.01 30.87 36.76 31.15 0.7511 1 1
DoRAr=16 28.25 26.26 28.53 33.38 29.00 0.7748 1 1
CoLA 36.28 32.25 40.88 46.44 38.63 0.9406 2 3
HydraLoRA 31.52 28.16 32.95 38.17 32.55 0.8267 1 3
MASA 35.96 32.70 42.31 46.93 39.05 0.8810 5 1
Table 8: Comparison of zero-shot performance (%) of our proposed MASA method against other PEFT techniques on the LLaMA3.2-3B base model. Performance is reported as accuracy on the MMLU benchmark, broken down into four domain subsets (Humanities, STEM, Social Sciences, Other) and their average. Columns AA and $B$ denote the number of $A$ and $B$ matrices in the respective LoRA architectures. The best accuracy in each column is highlighted in bold.
Baselines
@@
  • Prompt Tuning: Prompt Tuning is a parameter-efficient fine-tuning method that learns ”soft prompts” - continuous prompt embeddings that are prepended to the input of frozen language models. Unlike discrete text prompts, soft prompts are learned through backpropagation and can incorporate signals from labeled examples. The method becomes more competitive with scale, matching full fine-tuning performance as models exceed billions of parameters while using significantly fewer trainable parameters.
@@
  • P-Tuning: P-Tuning employs trainable continuous prompt embeddings to enhance the natural language understanding capabilities of GPT-style models. The method introduces learnable virtual tokens that are optimized during fine-tuning while keeping the pre-trained model parameters frozen. P-Tuning v2 extends this concept with deep prompt tuning, applying continuous prompts to every layer input of the transformer, significantly improving performance universally across scales and tasks.
@@
  • LoRA: Low-Rank Adaptation (LoRA) freezes pre-trained model weights and injects trainable rank decomposition matrices into each layer of the transformer architecture. The method hypothesizes that weight updates during adaptation have a low intrinsic rank, decomposing the update matrix into the product of two low-rank matrices. LoRA dramatically reduces trainable parameters while maintaining comparable performance to full fine-tuning.
@@
  • DoRA: Weight-Decomposed Low-Rank Adaptation (DoRA) decomposes pre-trained weights into magnitude and direction components for fine-tuning. It employs LoRA for directional updates while separately training the magnitude vector, effectively enhancing both learning capacity and training stability of LoRA. DoRA consistently outperforms LoRA across various downstream tasks while avoiding additional inference overhead.
@@
  • VB-LoRA: VB-LoRA introduces a “divide-and-share” paradigm that breaks the barriers of low-rank decomposition across matrix dimensions, modules, and layers by sharing parameters globally via a vector bank. The method composes all the low-rank matrices of LoRA from a shared vector bank with a differentiable top-k admixture module. VB-LoRA achieves extreme parameter efficiency while maintaining comparable or better performance compared to state-of-the-art PEFT methods, using only 0.4% of LoRA’s stored parameters when fine-tuning large language models.
@@
  • BSLoRA: Bi-Share LoRA (BSLoRA) extends local LoRA with intra-LoRA and inter-LoRA parameter sharing to better capture both local and global information simultaneously. The method reduces trainable parameters through both intra-layer and inter-layer parameter sharing while maintaining or even enhancing model performance. BSLoRA incorporates three transformation methods to improve the compatibility and collaborative efficiency of shared parameters with varying shapes, achieving superior performance with only 44.59% of the parameters of standard LoRA.
@@
  • CoLA: Collaborative Low-Rank Adaptation (CoLA) introduces a flexible LoRA architecture that decouples the number of low-rank matrices A and B, enabling collaborative strategies for multi-task learning. CoLA employs task-agnostic and task-specific modules with three collaborative strategies: fully collaborative, random collaborative, and heuristic collaborative approaches, effectively mitigating task interference while maintaining parameter efficiency.
@@
  • HydraLoRA: HydraLoRA is an asymmetric fine-tuning architecture that effectively identifies and adapts to intrinsic data components, such as sub-domains or diverse tasks. It allocates distinct B matrices for task-specific features, while a shared A matrix integrates global information, enabling efficient parameter utilization and enhanced performance. The architecture employs a trainable MoE router to automatically segregate training samples and dynamically merge multiple B matrices during inference.
@@
  • MultiLoRA: MultiLoRA addresses the limitation of LoRA’s explicit low-rank constraint in complex multi-task scenarios by reducing the dominance of the top singular vectors. The method scales LoRA modules horizontally and changes parameter initialization to reduce parameter dependency, yielding more balanced unitary subspaces. MultiLoRA democratizes LoRA adaptation for better multi-task learning by enabling more effective knowledge sharing across diverse tasks.
@@
  • LoRAMoE: LoRAMoE introduces several low-rank adapters (LoRA) and integrates them using a router network, functioning as a plugin version of Mixture of Experts (MoE). The method freezes the backbone model and forces a portion of LoRAs to focus on leveraging world knowledge to solve downstream tasks, effectively alleviating world knowledge forgetting during supervised fine-tuning. LoRAMoE addresses the challenge that large-scale increases in instruction data can damage the world knowledge previously stored in large language models.
@@
  • MTL-LoRA: Multi-Task Learning LoRA (MTL-LoRA) enhances LoRA’s multi-task learning capability by incorporating task-specific transformations in low-rank space along with adaptive exploration of multiple information sharing methods. The approach augments LoRA with additional task-adaptive parameters that differentiate task-specific information while capturing shared knowledge across various tasks within low-dimensional spaces, enabling effective joint adaptation to different target domains.
@@
Table 9: Experimental hyperparameter settings.
Hyperparameters
Appendix E Appendix D Additional Experimental Results
Expert Matrix Specialization Analysis
Cross-Model Adaptation Analysis
Table 8 complements the results in Section 4 by benchmarking all Parameter-Efficient Fine-Tuning (PEFT) baselines on the smaller 3 B-parameter LLaMA3.2-3B backbone. Consistent with the findings on LLaMA3-8B, our method MASA achieves the highest zero-shot accuracy on MMLU, reaching an average of 39.05 %. This represents a 0.42 pp gain over second best baseline, CoLA, while requiring 6.3 % fewer trainable parameters. MASA also secures column-leading results in three of the four domain subsets (STEM, Social Sciences, and Other) with a configuration of $A=5$ and $B=1$. Relative to LoRAand HydraLoRA, MASA delivers gains of 7.9 and 6.5 percentage points, respectively. Preservation of this performance across both the 8B and 3B backbones substantiates the scalability and robustness of MASA under varying model capacities and parameter budgets.

### v2

@@ Appendix A Rationale for Summation-Based Aggregation
Token-level routers, introduced by the Sparsely-Gated layer(Shazeer et al., 2017) and subsequently adopt in GShard (Lepikhin et al., 2020) and Switch-Transformer(Fedus et al., 2022), enable conditional computation (top-$k$ expert selection) and explicit expert specialization. These benefits explain their popularity in large-scale language and vision models.
@@
  • SmoothMoE (Xu et al., 2025) replaces top-$k$ gating with a soft mask. With identical expert backends, the dense variant stays within $0.1$ BLEU of the routed model on WMT En-De while reducing gradient variance and eliminating auxiliary losses.
  • On ImageNet, V-MoE (Riquelme et al., 2021), reports that a dense (all-expert) mixture narrows the accuracy gap with routed ViT-MoE to less than 0.30.3pp, while avoiding load-balancing issues.
@@
The proposed architecture prioritizes feature-side parameter enrichment through multi-$A$ experts while implementing asymmetric cross-layer sharing (ACS) to achieve parameter efficiency. The uniform summation mechanism, expressed as $\Delta W_{\text{MAE}}=B\left(\sum_{i=1}^{N}A_{i}\right)$, provides three key advantages:
@@
    Empirical sufficiency: While sparse routing strategies combined with effective load balancing have been proven to facilitate expert differentiation, our visualization results reveal that even a simple summation strategy can achieve effective expert specialization. The experimental results presented in Figure 2 provide clear evidence of task differentiation among our Expert A matrices, demonstrating that summation mechanism is sufficient for achieving functional specialization.
@@
Appendix B Additional Experimental Results
Table 7: Comparison of zero-shot performance (%) of our proposed MASA method against other PEFT techniques on the LLaMA3.2-3B base model (Grattafiori et al., 2024). Performance is reported as accuracy on the MMLU benchmark, broken down into four domain subsets (Humanities, STEM, Social Sciences, Other) and their average. Columns AA and $B$ denote the number of $A$ and $B$ matrices in the respective LoRA architectures. The best accuracy in each column is highlighted in bold.
Model MMLU Params. (%) AA $B$
Hum. STEM Social Other Avg.
LLaMA3.2-3B 24.36 21.15 21.90 23.98 23.02
Prompt Tuning 25.36 24.52 24.50 23.95 25.36 0.0008
P-Tuning 25.08 25.06 21.81 24.46 24.44 0.0530
LoRAr=8 26.14 24.26 25.58 28.81 26.19 0.3770 1 1
LoRAr=16 29.73 28.01 30.87 36.76 31.15 0.7511 1 1
DoRAr=16 28.25 26.26 28.53 33.38 29.00 0.7748 1 1
CoLA 36.28 32.25 40.88 46.44 38.63 0.9406 2 3
HydraLoRA 31.52 28.16 32.95 38.17 32.55 0.8267 1 3
MASA 35.96 32.70 42.31 46.93 39.05 0.8810 5 1
B.1 Additional Backbone Results
Table 7 complements the results in Section 4 by benchmarking all Parameter-Efficient Fine-Tuning (PEFT) baselines on the smaller 3 B-parameter LLaMA3.2-3B (Grattafiori et al., 2024) backbone. Consistent with the findings on LLaMA3-8B (Grattafiori et al., 2024), our method MASA achieves the highest zero-shot accuracy on MMLU (Hendrycks et al., 2020), reaching an average of 39.05 %. This represents a 0.42pp gain over second best baseline, CoLA (Zhou et al., 2025a), while requiring 6.3 % fewer trainable parameters. MASA also secures column-leading results in three of the four domain subsets (STEM, Social Sciences, and Other) with a configuration of $A=5$ and $B=1$. Relative to LoRA (Hu et al., 2022) and HydraLoRA (Tian et al., 2024), MASA delivers gains of 7.9 and 6.5 percentage points, respectively. Preservation of this performance across both the 8B and 3B backbones substantiates the scalability and robustness of MASA under varying model capacities and parameter budgets.
We further evaluate MASA on Qwen2.5-7B-Instruct and Mistral-7B; the per-category MMLU breakdowns are reported in Table 8. Due to baseline availability and compute budget, these additional-backbone runs compare MASA against the strongest applicable PEFT baselines (LoRA and HydraLoRA) under the same rank setting.
Table 8: Additional backbone results on MMLU. We report average MMLU and three subject groups. All methods use the same rank ($r=16$). (#A,#B)(\#A,\#B) indicates the numbers of down-projection matrices $A$ and up-projection matrices $B$.
Qwen2.5-7B-Instruct (MMLU)
Method Avg Law Medical Other (#A,#B)(\#A,\#B)
LoRA 71.71 55.30 73.79 74.10 (1,1)
HydraLoRA 71.47 54.74 72.70 72.03 (1,3)
MASA 71.91 55.53 74.40 74.24 (5,1)
Mistral-7B (MMLU)
Method Avg Law Medical Other (#A,#B)(\#A,\#B)
LoRA 58.83 47.93 64.03 59.90 (1,1)
HydraLoRA 58.18 47.93 63.21 59.16 (1,3)
MASA 58.99 48.33 64.71 59.96 (5,1)
B.2 Expert Matrix Specialization Analysis
Appendix C Experimental Setting
C.1 Datasets and Benchmarks
@@
Table 9: Overview of the datasets employed in this study.
Dataset Lang. Source / Version Task Types
[…]
C.2 Baselines
@@
  • Prompt Tuning: Prompt Tuning (Lester et al., 2021) is a parameter-efficient fine-tuning method that learns ”soft prompts” - continuous prompt embeddings that are prepended to the input of frozen language models. Unlike discrete text prompts, soft prompts are learned through backpropagation and can incorporate signals from labeled examples. The method becomes more competitive with scale, matching full fine-tuning performance as models exceed billions of parameters while using significantly fewer trainable parameters.
@@
  • P-Tuning: P-Tuning (Liu et al., 2021) employs trainable continuous prompt embeddings to enhance the natural language understanding capabilities of GPT-style models. The method introduces learnable virtual tokens that are optimized during fine-tuning while keeping the pre-trained model parameters frozen. P-Tuning v2 extends this concept with deep prompt tuning, applying continuous prompts to every layer input of the transformer, significantly improving performance universally across scales and tasks.
@@
  • LoRA: Low-Rank Adaptation (LoRA) (Hu et al., 2022) freezes pre-trained model weights and injects trainable rank decomposition matrices into each layer of the transformer (Vaswani et al., 2017) architecture. The method hypothesizes that weight updates during adaptation have a low intrinsic rank, decomposing the update matrix into the product of two low-rank matrices. LoRA dramatically reduces trainable parameters while maintaining comparable performance to full fine-tuning.
@@
  • DoRA: Weight-Decomposed Low-Rank Adaptation (DoRA) (Liu et al., 2024) decomposes pre-trained weights into magnitude and direction components for fine-tuning. It employs LoRA for directional updates while separately training the magnitude vector, effectively enhancing both learning capacity and training stability of LoRA. DoRA consistently outperforms LoRA across various downstream tasks while avoiding additional inference overhead.
@@
  • VB-LoRA: VB-LoRA (Li et al., 2024) introduces a “divide-and-share” paradigm that breaks the barriers of low-rank decomposition across matrix dimensions, modules, and layers by sharing parameters globally via a vector bank. The method composes all the low-rank matrices of LoRA from a shared vector bank with a differentiable top-k admixture module. VB-LoRA achieves extreme parameter efficiency while maintaining comparable or better performance compared to state-of-the-art PEFT methods, using only 0.4% of LoRA’s stored parameters when fine-tuning large language models.
@@
  • BSLoRA: Bi-Share LoRA (BSLoRA) (Zhou et al., 2025b) extends local LoRA with intra-LoRA and inter-LoRA parameter sharing to better capture both local and global information simultaneously. The method reduces trainable parameters through both intra-layer and inter-layer parameter sharing while maintaining or even enhancing model performance. BSLoRA incorporates three transformation methods to improve the compatibility and collaborative efficiency of shared parameters with varying shapes, achieving superior performance with only 44.59% of the parameters of standard LoRA.
@@
  • CoLA: Collaborative Low-Rank Adaptation (CoLA) (Zhou et al., 2025a) introduces a flexible LoRA architecture that decouples the number of low-rank matrices A and B, enabling collaborative strategies for multi-task learning. CoLA employs task-agnostic and task-specific modules with three collaborative strategies: fully collaborative, random collaborative, and heuristic collaborative approaches, effectively mitigating task interference while maintaining parameter efficiency.
@@
  • HydraLoRA: HydraLoRA (Tian et al., 2024) is an asymmetric fine-tuning architecture that effectively identifies and adapts to intrinsic data components, such as sub-domains or diverse tasks. It allocates distinct B matrices for task-specific features, while a shared A matrix integrates global information, enabling efficient parameter utilization and enhanced performance. The architecture employs a trainable MoE router to automatically segregate training samples and dynamically merge multiple B matrices during inference.
@@
  • MultiLoRA: MultiLoRA (Wang et al., 2023b) addresses the limitation of LoRA’s explicit low-rank constraint in complex multi-task scenarios by reducing the dominance of the top singular vectors. The method scales LoRA modules horizontally and changes parameter initialization to reduce parameter dependency, yielding more balanced unitary subspaces. MultiLoRA democratizes LoRA adaptation for better multi-task learning by enabling more effective knowledge sharing across diverse tasks.
@@
  • LoRAMoE: LoRAMoE (Dou et al., 2024) introduces several low-rank adapters (LoRA) and integrates them using a router network, functioning as a plugin version of Mixture of Experts (MoE). The method freezes the backbone model and forces a portion of LoRAs to focus on leveraging world knowledge to solve downstream tasks, effectively alleviating world knowledge forgetting during supervised fine-tuning. LoRAMoE addresses the challenge that large-scale increases in instruction data can damage the world knowledge previously stored in large language models.
@@
  • MTL-LoRA: Multi-Task Learning LoRA (MTL-LoRA) (Yang et al., 2025) enhances LoRA’s multi-task learning capability by incorporating task-specific transformations in low-rank space along with adaptive exploration of multiple information sharing methods. The approach augments LoRA with additional task-adaptive parameters that differentiate task-specific information while capturing shared knowledge across various tasks within low-dimensional spaces, enabling effective joint adaptation to different target domains.
C.3 Hyperparameters
Table 10: Experimental hyperparameter settings.